In a LAN, assume that the rate of transmission is 1 Mbps, the speed of propagation is 108 meter/second, and the length of the LAN is 100 meters. Answer the following questions:

a. How long does it take to send out a frame of 1000 bits (in miliseconds)?

b. How long does it take for the first bit to reach the end of the LAN (in milisecond)
Answer #1

Given,

     Bandwidth(bw) = 1Mbps

      Speed(v) = 108 m/s

      Lenght (d) = 100 m

a) They are asking Transmission time (Tt). that is equals to frame size/bandwidth.

       Tt = 1000 bits/1 Mbps = 103/106 sec = 10-3 sec = 1 msec

b) Propagation time (Tp) = length/speed

       Tp = 100m/108mps = 0.9259 sec = 9.25X10-4 msec